solve by substitution
3m + n =1
m - 4n = 9

Answer by LarissaRichardson(34)

3m+n=1
m-4n=9
Find what n is in the first equation (we find n because it doesn't have a coefficient and will be easier.)
3m+n=1
n=1-3m
Now that you have found what n is, substitute it in the second equation as:
m-4(1-3m)=9
Distribute now:
m-4+12m=9
Combine like terms:
13m=13
Divide by 13
m=1
Now to find n, substitute the m=1 into either equation. I will use the first equation:
3(1)+n=1
3+n=1
Combine like terms
n=-2
Now check your answers by plugging them in to the equations:
3m+n=1
3(1)+(-2)=1
3-2=1 YAY!
Now check your other equation!!